# FlakyFix: Using Large Language Models for Predicting Flaky Test Fix Categories and Test Code Repair

### v3

@@ Abstract
Flaky tests are problematic because they non-deterministically pass or fail for the same software version under test, causing confusion and wasting development effort. While machine learning models have been used to predict flakiness and its root causes, there is much less work on providing support to fix the problem. To address this gap, in this paper, we focus on predicting the type of fix that is required to remove flakiness and then repair the test code on that basis. We do this for a subset of flaky tests where the root cause of flakiness is in the test itself and not in the production code. One key idea is to guide the repair process with additional knowledge about the test’s flakiness in the form of its predicted fix category. Thus, we first propose a framework that automatically generates labeled datasets for 13 fix categories and trains models to predict the fix category of a flaky test by analyzing the test code only. Our experimental results using code models and few-shot learning show that we can correctly predict most of the fix categories. To show the usefulness of such fix category labels for automatically repairing flakiness, we augment the prompts of GPT 3.5 Turbo, a Large Language Model (LLM), with such extra knowledge to request repair suggestions. The results show that our suggested fix category labels, complemented with in-context learning, significantly enhance the capability of GPT 3.5 Turbo in generating fixes for flaky tests. Based on the execution and analysis of a sample of GPT-repaired flaky tests, we estimate that a large percentage of such repairs, (roughly between 70% and 90%) can be expected to pass. For the failing repaired tests, on average, 16% of the test code needs to be further changed for them to pass.
@@ 1 Introduction
Our findings indicate several key points: (1) Smaller pre-trained code models demonstrate a high accuracy in predicting most fix categories, achieving precision rates ranging from 83% to 98% and recall rates from 67% to 97%, solely by analyzing the test code; (2) UniXcoder surpasses CodeBERT in predicting fix categories; (3) The inclusion of FSL does not yield significant improvements in the results; (4) Our predicted fix category labels play a significant role in guiding GPT to generate better full fixes for flaky tests. We evaluated this using CodeBLEU [29] and achieved a mean score of 81% when using predicted fix category labels in the input prompt, compared to 75% without them. These results are further enhanced with in-context learning [30], which we tested on three different fix categories; (5) Based on the execution and analysis of a sample of GPT-repaired tests, we estimate that a large percentage can be expected to pass, roughly between 70% and 90% of them. For the generated tests that fail, the required changes to correct them are limited, with an average of 16% of the tokens needing modification in the test code. In summary, the contributions of this paper are as follows:
@@
  • Based on the execution of a sample of GPT-generated tests, we use CodeBLEU metrics combined with Bootstrapping statistics to estimate the pass rate for all generated repaired tests, including those that could not be executed. We also assess the changes required to manually fix tests that failed.
@@ 2.2.2 Fine-tuning Code Models
As discussed earlier, we have fine-tuned CodeBERT and UniXcoder independently with two different techniques. Given our limited labeled training data, which is a common situation in our application context, few shot learning approaches [27], such as the Siamese Network architecture [44], are adequate candidate solutions. Therefore, similar to existing work [23], We constructed a Siamese Network [45, 46, 23] with three identical sub-networks. Each sub-network is based on either CodeBERT or UniXcoder, essentially using the same model (weights and network parameters) three times within the network structure. The inputs to this Siamese Network are anchor, positive and negative tests, which are processed by each sub-network of the Siamese Network in parallel.
Given a category (e.g., Change Assertion), the anchor is a flaky test in the training dataset that belongs to the positive class of that category, e.g., it is labeled as Change Assertion. A positive sample is another flaky test in the training dataset that belongs to the same class as the anchor (e.g., labeled again as Change Assertion) and is considered similar to the anchor. The negative sample is another flaky test from the dataset that belongs to a different class than that of the anchor (e.g., NOT labeled as Change Assertion) and is considered dissimilar to the anchor. More details about the procedure to create these samples in our experiment will be provided in Section 3.
By training the Siamese Network with a specific loss function (Triplet Loss Function [47]), the model learns to generate [CLS] embeddings for each anchor test case in a way that it captures the similarity between the anchor and positive examples while minimizing the similarity between the anchor and negative examples.
@@ 3.2.1 Existing Datasets
The largest publicly available and relevant dataset is the International Dataset of Flaky Tests (IDoFT) [^13] https://mir.cs.illinois.edu/flakytests, an open-source dataset that is continuously updated. In this research, we collected a Java dataset of flaky tests from IDoFT until October 23, 2023. This dataset comprises 5500 tests, including 128 duplicate tests. After excluding duplicates, the remaining 5372 tests were analyzed, revealing that 562 tests were fixed solely by modifying the test code and are therefore aligned with our research objectives. As explained in the Introduction section as well, the flaky tests that can be fixed in a black-box manner are in the minority, but they are still important since testers and testing tools can fix them even without any access to production code, which is typical for some QA teams. We ensured that these fixes were accepted, meaning that the pull requests aimed at resolving these flaky tests on GitHub were approved by the original developers. The remaining tests were resolved through test configuration adjustments changes in test case execution order, fixes implemented in the production code, or are pending in open pull requests (awaiting approval, rejection, or being deliberately skipped by the original developers). More details about this can be found online on the IDoft Repository.
Consequently, our final dataset comprises 562 flaky tests and their corresponding resolutions from the IDoFT Java dataset. Notably, this dataset represents the largest available repository of flaky tests that are fixable through modifications of the test code. These tests originate from 123 distinct projects, offering a broad spectrum of test cases, which greatly benefits the generalizability of our predictive models.
@@ 3.2.2 Labeled Dataset of Flaky Test Fix Categories
To create the labeled dataset out of the IDoFT dataset, we followed the procedure discussed in Section 2 and labeled each flaky test automatically. Our dataset is publicly available in our replication package  [^14] This will be available online, once this paper is accepted.. The resulting labeled dataset is unbalanced with respect to the distribution of tests across fix categories. Though, in our prediction phase, to handle the lack of data, we apply Few-Shot learning, to effectively predict the fix categories will require a minimum number of test cases per category. Therefore, we selected the categories with at least 30 tests as our training set. As a result, our labeled dataset consists of the top nine fix categories outlined in Table II for training our model. Since a flaky test may be associated with multiple fix categories, we have transformed our multi-label dataset into nine distinct bi-labeled sets, one per category. In each of the nine sets, the positive class includes tests that belong to the corresponding fix category and the negative class includes all the other tests. This allows us to treat the problem of predicting each label (category) as a separate binary classification problem.
@@ 3.3.3 Prompt Design
To get the fixed flaky test code from GPT, we use three different types of prompts, as follows: (1) we pass the flaky test code in the prompt to the model and ask it to generate its fix as shown in Figure 6, (2) we provide both the flaky test code and the fix category label we predicted for that flaky test. We then ask the GPT model to generate a fix as shown in Figure 7, and (3) we extend the content of the second type by also passing a few examples to the model as shown in Figure 8. This type of prompting aims to assess GPT’s performance with in-context learning. We focused on three fix categories comprising more than 30 instances (to have large enough sample sizes to make meaningful conclusions) within our dataset of 181 tests: Change Assertion (64 examples), Change Data Structure (41 examples), and Change Condition (41 examples).
For in-context learning, from each of these three fix categories, we randomly selected five flaky tests, along with their respective fixes and fix category labels, as prompts for the GPT model. Our prompt requested the model to provide a fix for a new flaky test based on a given fix category label, utilizing information derived from these example tests and their associated fix category labels. Our prompt explicitly requests the model to provide only the fixed test code, excluding any additional textual descriptions. This approach streamlines the output, ensuring that the model generates the test code directly without requiring post-processing. In the future, we plan to further test our GPT model using a larger number of examples of flaky tests. This expansion aims to evaluate and enhance the model’s capabilities in handling varied scenarios.
@@ 3.3.5 Evaluation Metrics
Beside syntactic similarity metrics like CodeBLEU, we also executed a subset of repaired tests to assess the correctness of GPT-repaired flaky tests. Analysing the execution results, we calculated the passing rate of the repaired tests. Then using bootstrapping [77] with a 95% confidence level, a statistical method for estimating properties of a large sample from repeated sampling (with replacement) from a small sample, we established lower and upper bound percentages of passing test cases within the entire flaky test set, including non-executable test cases. Specifically, using these bounds, we estimated the passing rate of remaining tests with CodeBLEU scores above or equal to 80%.
@@ 3.4.3 RQ3 Results
Now to expand the analysis to the third prompting approach (in-context learning), Table VI compares the three approaches for the Change Assertion, Change Data Structure, and Change Condition fix categories. As explained, we only look at these categories since after removing 5 samples per category to be included in the context, only these categories have enough samples to do a proper analysis (run statistical tests).
@@
Looking at the table, we see that, for the Change Data Structure category, the mean and median CodeBLEU scores increase from 80% to 86% and 87% to 94%, respectively, when we use in-context learning versus simply providing the labels. Similarly, for the  Change Condition category, mean and median CodeBLEU scores increase from 88% to 92% and 90% to 97%, respectively. However, in the case of the Change Assertion category, there is no discernible improvement in the model’s performance after exposure to a limited number of examples. This could be attributed to the variability of assertion statements within the test code, suggesting that multiple approaches may exist to fix assertions and resolve flakiness. It is plausible that the number and variety of examples provided was insufficient for the model to grasp the diverse spectrum of assertion-related fixes. Further, as depicted in the boxplots in Figure  9, the overall distribution of the CodeBLEU scores shows significant variance for all types of prompts and the three fix categories. The low minimum values in all cases shows that there are flaky tests that are not fixed regardless of the prompting strategy. However, when looking at the CodeBLEU over all three categories for in-context learning, it is clear that, for the majority of the cases, the CodeBLEU scores are quite high. For example, even for Change Assertion category, which has the lowest minimum scores, 75% of samples have CodeBLEUs higher than 55%, and for half of the test cases the CodeBLEU score are in fact greater than 85%.
@@
Given the variance in CodeBLEU values per category, we again used the non-parametric Wilcoxon signed-rank test with a significance level of $\alpha=0.05$ to make sure the improvements we see in medians are statistically significant as well. Our findings indicate that when accounting for all three categories, both for prompts with label and without label, in-context learning makes a significant difference, with p-values below the significance threshold (0.00024 and 0.0000001, respectively). However, looking at individual categories, this difference for prompts with label is only significant for  Change Data Structure and  Change Condition, with p-values of 0.00024 and 0.0072, respectively. In future investigations, we intend to systematically explore the use of different prompts by augmenting the number and type of examples or experimenting with other prompt engineering strategies [74]. Further details regarding GPT’s performance on individual fix categories are discussed in Section 4.
@@
To gain deeper insights into the performance of GPT-repaired flaky tests, we conducted a series of evaluations based on test execution results. For the reasons explained in Section 3.2.5, we could only execute a total of 35 tests, each repeated 10 times, resulting in 24 passing and 11 failing tests (all with consistent results across the 10 runs). Our analysis focused on correlating the execution outcomes with the corresponding CodeBLEU scores to ascertain the reliability of the CodeBLEU metric in assessing the generated test. Specifically, we expected tests with higher CodeBLEU scores to have a higher likelihood of passing and vice-versa. Among the passing tests, 16 achieved a perfect CodeBLEU score of 100%, indicating an exact match with the developer-repaired versions. Additionally, 22 tests scored above or equal to 80% on CodeBLEU. However, one passing test exhibited a notably lower CodeBLEU score of 36%, as illustrated in Figure 10, showing an alternative but correct repair.
Conversely, among the failed tests, five achieved CodeBLEU scores above or equal to 80%. Utilizing the subset of 27 tests with CodeBLEU scores of 80% or higher (comprising 22 passing and 5 failed tests), we employed bootstrapping to establish a 95% confidence interval for percentages of passing test cases, thus obtaining estimates of 63% (lower bound) and 93% (upper bound). Based on these bounds, we then estimated the passing rates for two datasets: one consisting of all 181 tests covering all fix categories, and another with 125 tests encompassing the three fix categories supplemented by in-context learning. Tables VII and VIII present the passing test estimates in terms of upper and lower bounds across different prompt settings, for the 181 and 125 tests, respectively. According to our estimates, within the 181-test dataset:
@@
  • With No Label prompt: 65 tests (68%) to 85 tests (89%) are projected to pass.
@@
  • With Label prompt: 81 tests (67%) to 108 tests (90%) are projected to pass.
In the 125-test dataset:
@@
  • With No Label prompt: 50 tests (70%) to 63 tests (89%) are projected to pass.
@@
  • With Label prompt: 60 tests (68%) to 79 tests (90%) are projected to pass.
@@
  • With in-context learning: 62 tests (69%) to 80 tests (98%) are projected to pass.
To conclude, the results suggest that a large percentage of test cases are passing, roughly somewhere in the 70-90% range. Note that the above estimates also include the 35 tests that were actually executed, along with the ones whose passing behavior was predicted using Bootstrapping results.
Furthermore, for the 11 failed tests, we calculated the edit distance, as a surrogate measure to assess the manual fixing effort. It revealed an average token replacement percentage of 16% to convert a failed test to the corresponding ground truth. Figure 11 depicts the trend between CodeBLEU scores and the percentage of token replacement. The scatter plot clearly suggests that GPT-repaired tests with higher CodeBLEU scores tend to require fewer edits for repair, suggesting close alignment with the developer-repaired versions. As mentioned earlier, all generated tests, along with their CodeBLEU scores and execution results, are provided in our replication package.
@@
RQ3 summary: In all fix categories, either providing category labels or applying in-context learning significantly boosts the GPT model’s ability to repair flaky tests. For two categories, in-context learning provides significantly better results than prompts with labels.  Furthermore, providing labels in prompts or with in-context learning, we project that a large percentage of repaired tests will pass, from 67% to 90% , and from 69% to 98%, for the two types of prompts respectively. For the tests that fail, required changes are rather limited, with an average of 16% of the tokens to be changed in test code.
@@
Prompts Lower Bound Upper Bound Number of Tests with CodeBLEU >=80%
@@
With No Label 65 68% 85 89% 95
With Label 81 67% 108 90% 120
TABLE VII: Passing Estimates (for all 181 test cases) for the GPT-repaired tests with and without providing the fix category labels in the GPT prompt.
@@
With No Label 50 70 % 63 89% 71
With Label 60 68% 79 90% 88
In-Context Learning 62 69% 88 98% 90
TABLE VIII: Passing Estimates for the GPT-repaired tests using fix category labels and in-context learning (for 125 test cases).

### v4

@@ Abstract
Flaky tests are problematic because they non-deterministically pass or fail for the same software version under test, causing confusion and wasting development effort. While machine learning models have been used to predict flakiness and its root causes, there is much less work on providing support to fix the problem. To address this gap, in this paper, we focus on predicting the type of fix that is required to remove flakiness and then repair the test code on that basis. We do this for a subset of flaky tests where the root cause of flakiness is in the test itself and not in the production code. One key idea is to guide the repair process with additional knowledge about the test’s flakiness in the form of its predicted fix category. Thus, we first propose a framework that automatically generates labeled datasets for 13 fix categories and trains models to predict the fix category of a flaky test by analyzing the test code only. Our experimental results using code models and few-shot learning show that we can correctly predict most of the fix categories. To show the usefulness of such fix category labels for automatically repairing flakiness, we augment the prompts of GPT 3.5 Turbo, a Large Language Model (LLM), with such extra knowledge to request repair suggestions. The results show that our suggested fix category labels, complemented with in-context learning, significantly enhance the capability of GPT 3.5 Turbo in generating fixes for flaky tests. Based on the execution and analysis of a sample of GPT-repaired flaky tests, we estimate that a large percentage of such repairs, (roughly between 51% and 83%) can be expected to pass. For the failing repaired tests, on average, 16% of the test code needs to be further changed for them to pass.
@@ 1 Introduction
Our findings indicate several key points: (1) Smaller pre-trained code models demonstrate a high accuracy in predicting most fix categories, achieving precision rates ranging from 83% to 98% and recall rates from 67% to 97%, solely by analyzing the test code; (2) UniXcoder surpasses CodeBERT in predicting fix categories; (3) The inclusion of FSL does not yield significant improvements in the results; (4) Our predicted fix category labels play a significant role in guiding GPT to generate better full fixes for flaky tests. We evaluated this using CodeBLEU [29] and achieved a mean score of 81% when using predicted fix category labels in the input prompt, compared to 75% without them. These results are further enhanced with in-context learning [30], which we tested on three different fix categories; (5) Based on the execution and analysis of a sample of GPT-repaired tests, we estimate that a large percentage can be expected to pass, roughly between 46% and 80% of them. For the generated tests that fail, the required changes to correct them are limited, with an average of 16% of the tokens needing modification in the test code. In summary, the contributions of this paper are as follows:
@@
  • Based on the execution of a sample of GPT-generated tests, we rely on CodeBLEU metrics and statistical methods, including bootstrapping and logistic regression, to estimate the pass rate for all generated repaired tests, including those that could not be executed. We also assess the magnitude of the modifications needed to manually fix tests that failed.
@@ 2.2.2 Fine-tuning Code Models
As discussed earlier, we have fine-tuned CodeBERT and UniXcoder independently with two different techniques. Given our limited labeled training data, which is a common situation in our application context, few shot learning approaches [27], such as the Siamese Network architecture [44], are adequate candidate solutions. A Siamese Network trained with a specific loss function (Triplet Loss Function [45]) is designed to determine the similarity between inputs by learning an embedding space [46]—a high-dimensional vector space where inputs are represented as embeddings or vectors. In this space, similar inputs are positioned closer together, while dissimilar inputs are placed farther apart. Similar to existing work [23], we constructed a Siamese Network [47, 48, 23] with three identical sub-networks. Each sub-network is based on either CodeBERT or UniXcoder, essentially using the same model (weights and network parameters) three times within the network structure. The inputs to this Siamese Network are anchor, positive and negative tests, which are transformed by each sub-network of the Siamese Network into high-dimensional embeddings [46]. The anchor and positive tests, which belong to the same class, are projected into the embedding space such that their distance is minimized, while the distance between the anchor and the negative test, belonging to a different class, is maximized. This process ensures that similar inputs are grouped together in the vector space, facilitating tasks such as similarity comparison or classification based on learned similarities. During training, the triplet loss function fine tunes the embeddings by computing similarity scores, encouraging the network to improve its ability to distinguish flaky tests belonging to different categories. In our approach, given a category (e.g., Change Assertion), the anchor is a flaky test in the training dataset that belongs to the positive class of that category, e.g., it is labeled as Change Assertion. A positive sample is another flaky test in the training dataset that belongs to the same class as the anchor (e.g., labeled again as Change Assertion) and is considered similar to the anchor. The negative sample is another flaky test from the dataset that belongs to a different class than that of the anchor (e.g., NOT labeled as Change Assertion) and is considered dissimilar to the anchor. More details about the procedure to create these samples in our experiment will be provided in Section 3.
@@ 3.2.1 Existing Datasets
The largest publicly available and relevant dataset is the International Dataset of Flaky Tests (IDoFT) [^13] https://mir.cs.illinois.edu/flakytests, an open-source dataset that is continuously updated. In this research, we collected a Java dataset of flaky tests from IDoFT until October 23, 2023. The dataset originally contained 5,500 tests, including 128 duplicates. After removing duplicates, we were left with 5,372 tests. We then filtered this set to include only those flaky tests for which the fixes were accepted, meaning that the pull requests (PRs) addressing these flaky tests on GitHub were approved by the original developers, thus resulting in 788 tests. From this subset of flaky tests, we analyzed the diff of the original flaky tests and their developer repairs. We removed 226 tests that had repairs exclusively in the code under test (CUT) with no changes to the test code, leaving 562 tests that were fixed by modifying the test code. As explained in the Introduction section as well, the flaky tests that can be fixed in a black-box manner are in the minority, but they are still important since testers and testing tools can fix them even without any access to production code, which is typical for some QA teams. The remaining tests were resolved through test configuration adjustments, changes in test case execution order, fixes implemented in the production code, or are pending in open pull requests (awaiting approval, rejection, or being deliberately skipped by the original developers). More details about this can be found online on the IDoft Repository. Consequently, our final dataset comprises 562 flaky tests and their corresponding resolutions from the IDoFT Java dataset. Notably, this dataset represents the largest available repository of flaky tests that are fixable through modifications of the test code. These tests originate from 123 distinct projects, offering a broad spectrum of test cases, which greatly benefits the generalizability of our predictive models.
@@ 3.2.2 Labeled Dataset of Flaky Test Fix Categories
To create the labeled dataset out of the IDoFT dataset, we followed the procedure discussed in Section 2 and labeled each flaky test automatically. Our dataset is publicly available in our replication package  [^14] https://figshare.com/s/47f0fb6207ac3f9e2351. The resulting labeled dataset is unbalanced with respect to the distribution of tests across fix categories. Though, in our prediction phase, to handle the lack of data, we apply Few-Shot learning, to effectively predict the fix categories will require a minimum number of test cases per category. Therefore, we selected the categories with at least 30 tests as our training set. As a result, our labeled dataset consists of the top nine fix categories outlined in Table II for training our model. Since a flaky test may be associated with multiple fix categories, we have transformed our multi-label dataset into nine distinct bi-labeled sets, one per category. In each of the nine sets, the positive class includes tests that belong to the corresponding fix category and the negative class includes all the other tests. This allows us to treat the problem of predicting each label (category) as a separate binary classification problem.
@@ 3.3.3 Prompt Design
To get the fixed flaky test code from GPT, we use three different types of prompts, as follows: (1) we pass the flaky test code in the prompt to the model and ask it to generate its fix as shown in Figure 6, (2) we provide both the flaky test code and the fix category label we predicted for that flaky test. We then ask the GPT model to generate a fix as shown in Figure 7, and (3) we extend the content of the second type by also passing a few examples to the model as shown in Figure 8. This type of prompting aims to assess GPT’s performance with in-context learning. We focused on three fix categories comprising more than 30 instances (to have large enough sample sizes to make meaningful conclusions) within our dataset of 181 tests: Change Assertion (64 examples), Change Data Structure (41 examples), and Change Condition (41 examples). This resulted in a dataset of 146 tests. From these, we randomly selected 5 examples from each category for in-context learning, leaving a total of 131 tests for our in-context learning experiments.
We used these selected examples of flaky tests, along with their respective fixes and fix category labels, as prompts for the GPT model. Our prompt requested the model to provide a fix for a new flaky test based on a given fix category label, utilizing information derived from these example tests and their associated fix category labels. Our prompt explicitly requests the model to provide only the fixed test code, excluding any additional textual descriptions. This approach streamlines the output, ensuring that the model generates the test code directly without requiring post-processing. In the future, we plan to further test our GPT model using a larger number of examples of flaky tests. This expansion aims to evaluate and enhance the model’s capabilities in handling varied scenarios.
@@ 3.3.5 Evaluation Metrics
Beside syntactic similarity metrics like CodeBLEU, we also executed a subset of repaired tests to assess the correctness of GPT-repaired flaky tests. Analysing the execution results, we calculated the passing rate of the repaired tests. Then using bootstrapping [78] with a 95% confidence level, a statistical method for estimating properties of a large sample from repeated sampling (with replacement) from a small sample, we established lower and upper bound percentages of passing test cases within the entire flaky test set, including non-executable test cases. Additionally, we used a logistic regression [79] model to predict the pass rates of non-executable tests based on their CodeBLEU scores in various prompt settings.
@@ 3.4.3 RQ3 Results
Now to expand the analysis to the third prompting approach (in-context learning), Table VI compares the three approaches for the Change Assertion, Change Data Structure, and Change Condition fix categories. As explained, we only look at these categories since after removing 5 samples per category to be included in the context, only these categories have enough samples to do a proper analysis and run statistical tests.
@@
Looking at the table, we see that, for the Change Data Structure category, the mean and median CodeBLEU scores increase from 80% to 86% and 87% to 94%, respectively, when we use in-context learning versus simply providing the labels. Similarly, for the  Change Condition category, mean and median CodeBLEU scores increase from 88% to 92% and 90% to 97%, respectively. However, in the case of the Change Assertion category, there is no discernible improvement in the model’s performance after exposure to a limited number of examples. This could be attributed to the variability of assertion statements within the test code, suggesting that multiple approaches may exist to fix assertions and resolve flakiness. It is plausible that the number and variety of examples provided was insufficient for the model to grasp the diverse spectrum of assertion-related fixes. Further, as depicted in the boxplots in Figure  11, the overall distribution of the CodeBLEU scores shows significant variance for all types of prompts and the three fix categories. The low minimum values in all cases shows that there are flaky tests that are not fixed regardless of the prompting strategy. However, when looking at the CodeBLEU over all three categories for in-context learning, it is clear that, for the majority of the cases, the CodeBLEU scores are quite high. For example, even for the Change Assertion category, which has the lowest minimum scores, 75% of samples have CodeBLEUs higher than 55%, and for half of the test cases the CodeBLEU scores are greater than 85%.
@@
Given the variance in CodeBLEU values per category, we again used the non-parametric Wilcoxon signed-rank test with a significance level of $\alpha=0.05$ to make sure the improvements we see in medians are statistically significant as well. Our findings indicate that, when accounting for all three categories, both for prompts with label and without label, in-context learning makes a significant difference, with p-values below the significance threshold $p<0.001$. However, looking at individual categories, this difference for prompts with label is only significant for Change Data Structure and Change Condition, with p-values $p<0.001$ and $p=0.007$, respectively.
In future investigations, we intend to systematically explore the use of different prompts by augmenting the number and types of examples or experimenting with other prompt engineering strategies [75]. Further details regarding GPT’s performance on individual fix categories are discussed in Section 4.
@@
To gain deeper insights into the performance of GPT-repaired flaky tests, we conducted a series of evaluations based on test execution results. For the reasons explained in Section 3.2.5, we could only execute a total of 35 tests, each repeated 10 times, resulting in 24 passing and 11 failing tests, with consistent results across the 10 runs.
From the 35 executable tests, generated with labels and examples, for the three categories where the latter are available, we would like to first estimate the passing rate of fixed tests while accounting for the uncertainty associated with such a limited sample. Using Bootstrapping, we estimate a 95% confidence interval for the percentage of passing tests, that ranges from 51% to 83%. This interval is informative for testers, who do not have access to CodeBLEU scores in practical testing scenarios, and would like to know what to expect from FlakyFix. Based on these results, we conclude that FlakyFix provides an effective solution in a large proportion of cases.
Next we are interested in determining the relationship between execution outcomes and the corresponding CodeBLEU scores to ascertain the reliability of the CodeBLEU metric in assessing the generated test. Specifically, we expect that tests with higher CodeBLEU scores to have a higher likelihood of passing and vice-versa. Overall, the average score among passing and failing tests is 94% and 74%, respectively, suggesting that higher CodeBLEU scores are an indicator of passing tests. In particular, among the passing tests, 16 out of 24 achieved a perfect CodeBLEU score of 100%, indicating an exact match with the developer-repaired versions. Additionally, 22 tests scored above or equal to 80% on CodeBLEU. Conversely, among the 11 failed tests, only five tests achieved CodeBLEU scores above or equal to 80%. However, one passing test exhibited a notably lower CodeBLEU score of 36%, as illustrated in Figure 12, showing an alternative but correct repair. This particular case perfectly illustrates the limitations of using CodeBLEU scores.
To quantitatively model the relationship between CodeBLEU scores and the test passing probabilities, we train a Logistic Regression (LR) [79] model on the 35 executable tests, using their pass/fail labels, achieving an accuracy of 80% (correctly predicted test cases). The choice of LR is due to the small sample and the fact that we have one explanatory variable, CodeBLEU, that is expected to have a monotonic relationship with the test passing probability. The model’s goodness of fit, measured by Chi-Square test [82], yields a p-value of 0.010, which indicates a statistically significant relationship between CodeBLEU scores and the probability for a test to pass, despite the small size of the sample. The fitted LR curve between CodeBLEU scores and the probability for a test to pass is shown in Figure 9, showing a sharp increase in passing probability beyond CodeBLEU scores of 50.
To estimate the passing test rate among the non-executable tests, we apply this trained LR model to predict the outcomes of these tests. We calculate 95% confidence intervals for the predicted probabilities using standard errors derived via the delta method [83]. This method allows us to approximate the variance of the predicted probabilities by first calculating the standard errors on the logit scale and then transforming these to the probability scale. With this, we ensure that the confidence intervals account for the non-linear relationship between the logit and the predicted probability, giving a reliable measure of uncertainty around our predictions. According to our estimates, from the 181 test dataset, we obtain the following probability confidence intervals and passing test estimates:
@@
  • With No Label prompt: Confidence Interval: 0.44–0.50 — 80 (44%) tests to 91 (50%) tests are projected to pass.
@@
  • With Label prompt: Confidence Interval: 0.51–0.57 — 92 (51%) tests to 103 (57%) tests are projected to pass.
From the 131-test dataset (those that we can apply in-context learning on), we obtain the following results:
@@
  • With No Label prompt: Confidence Interval: 0.44–0.52 — 58 (44%) tests to 68 (52%) tests are projected to pass.
@@
  • With Label prompt: Confidence Interval: 0.50–0.58 — 66 (50%) tests to 76 (58%) tests are projected to pass.
@@
  • With in-context learning: Confidence Interval: 0.53–0.60 — 69 (53%) tests to 79 (60%) tests are projected to pass.
Tables VII and VIII provide the passing test estimates in terms of upper and lower bounds across different prompt settings, for the 181 and 131 test datasets, respectively.
From the above results, we can conclude that providing labels with the prompt to GPT does indeed make a difference in terms of passing tests. Overall, a reasonable proportion of generated tests with labels are predicted to pass, ranging approximately from 50% to 60% when accounting for uncertainty. Furthermore, for the 11 GPT-generated tests that failed upon execution, we calculated the edit distance as a surrogate measure to assess the manual effort required for fixing. The results revealed that, on average, 16% of the tokens needed to be replaced to convert a failed test into the corresponding ground truth.Figure 10 depicts the trend between CodeBLEU scores and the percentage of token replacement. The scatter plot suggests that GPT-repaired tests with higher CodeBLEU scores tend to require fewer edits for repair, suggesting closer alignment with the developer-repaired versions. Lastly, to gain further insights into GPT’s performance, we analyzed the fix categories of the 11 failed tests. Among these, six tests fell under the Change Data Format category. Of the remaining five, three were categorized under Change Assertion and two under Change Data Structure. The predominance of failed tests in the Change Data Format category can be attributed to the absence of in-context learning for this category, due to insufficient available instances as explained in Section 3.3.3. Our approach for this category relied solely on prompts with label information, affecting the quality of generations. However, as visible in Table VI, for the other three fix categories, the results indicate notable improvements in GPT’s generation quality with the implementation of in-context learning. As mentioned earlier, all generated tests, along with their CodeBLEU scores and execution results, are provided in our replication package.
@@
RQ3 summary: In all fix categories, either providing category labels or applying in-context learning significantly boosts the GPT model’s ability to repair flaky tests. For two categories, where in-context learning is possible, it provides better results than prompts with labels.  Furthermore, when providing labels in prompts and with in-context learning, when the latter is possible, we project that a large percentage $(>50\%)$ of repaired tests will pass. For the tests that fail, required changes are rather limited, with an average of 16% of the tokens to be changed in test code.
Figure 9: The Logistic Regression Curve showing a relation between CodeBLEU scores and probability of passing.
@@
Prompts Lower Bound Upper Bound
@@
With No Label 80 44% 91 50%
With Label 92 51% 103 57%
TABLE VII: Passing Estimates from the 181 test dataset for the GPT-repaired tests with and without providing the fix category labels in the GPT prompt.
@@
With No Label 58 44% 68 52%
With Label 66 50% 76 58%
In-Context Learning 69 53% 79 60%
TABLE VIII: Passing Estimates for the GPT-repaired tests using fix category labels and in-context learning (for the 131 tests).